OpenStudy (kobeni-chan):

Solve the radical equation: √(2x+3)=x (will medal) The choices are: x=3 and x=-1 x=3 x=-1 No Solution I worked it out and at the end of the quadratic equation, I got (x^2)-2x-3=0, with (x-3)(x+1) I don't understand how they got the positive and negative signs the other way around in the choices. I know the answer is x=3, but I think I just made a small mistake somewhere?

OpenStudy (aum):

x^2 - 2x - 3 = 0 (x-3)(x+1) = 0 x - 3 = 0 or x + 1 = 0 x = 3 or x = -1 But we have to test each solution with the original problem to make sure they are valid solutions. Put x = 3 in the original equation: √(2x+3)=x x = 3 √(6+3)=3 3 = 3 √(2x+3)=x x = -1 √(-2+3)=-1 1 is not equal to -1 So x = -1 is extraneous solution. x = 3

OpenStudy (kobeni-chan):

Ohh ok, I forgot to do that step in the beginning that switched the signs around. Thanks @aum

OpenStudy (aum):

You are welcome. If (x+a) is a factor then x = -a is a root / zero / solution. If (x-a) is a factor then x = a is a root / zero / solution.